OpenStudy (anonymous):

test for symmetry for: 11x^4 - 2y^3 = 10

OpenStudy (anonymous):

does anyone know this?

OpenStudy (anonymous):

You have to test for symmetry about every axis. The x, the y, and the origin. So testing for the x axis origin, set x = to -x. If the equation is the same, then it is symmetric about the x. In any case, this should only happen in even powers...regardless, (-11x^4)-2y^3=10 is still (11x^4)-2y^3=10. Do the same for y. You have an odd power, so (-2y^3) is not the same as 2y^3 and there is no symmetry about the y axis. For origin symmetry, set them both equal to negative values. Because the equation comes out to have x, but not y, symmetry, it is not symmetrical about the origin either.